OpenStudy (anonymous):

ln 4x=2.1 solve the log. equation algebraically. round the result to three decimal places

OpenStudy (anonymous):

ln 4x = 2 4x = e^2 x = (e^2)/4

OpenStudy (anonymous):

\[\ln(4x)=2.1\iff 4x=e^{2.1}\] etc